# Pluriclosed metrics on compact semisimple Lie groups

### v2

@@ Abstract.
Given a compact semisimple Lie group $G$ and a maximal torus $T\subset G$, we give an explicit description of all left and $\operatorname{Ad}(T)$-invariant pluriclosed Hermitian structures on $G$ in terms of the corresponding root system. They depend on $12d+$ 1 parameters in the irreducible case, where $\dim{T}=2d$. As applications, we obtain that the only left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics which are also CYT are bi-invariant metrics (i.e., Bismut flat) and study the pluriclosed flow as a neat ODE system.
@@ 1. Introduction
As far as we know, beyond $\mathrm{SU}(2)\times\mathrm{SU}(2)$, $\mathrm{SU}(3)$ and $\mathrm{SO}(9)$, the only other known examples of pluriclosed left-invariant metrics (non bi-invariant) on compact semisimple Lie groups were exhibited in [Ph] for $G=\mathrm{G}_{2}$.
@@
If $\operatorname{rk}(G)=2d$, then each of these metrics is compatible with a $d(d-1)$-parametric space of left-invariant complex structures on $G$. According to the theorem, for a fixed $J$, the space of all left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics compatible with $J$ depends on $r(2d+1)$ parameters and contains an $r$-parametric space of bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold. In particular, any irreducible $J$ admits a unique compatible bi-invariant metric up to scaling. Note that $J$ can be irreducible on a non-simple $G$ (see Example 2.1).
@@
and $F$ is the functional defined in (1). This converts $F$ in a Lyapunov function for the dynamical system in the sense that $F(X(t))$ is strictly decreasing along any solution $X(t)$, which implies that any $X(t)$ converges to $(1,\dots,1)$ as $t\to\infty$. Any left and $\operatorname{Ad}(T)$-invariant pluriclosed solution $g(t)$ therefore converges, as $t\to\infty$, to a bi-invariant metric (see Theorem 5.1 for a more precise statement). We note that the ODE systems (2) and (3) allows us to actually see the pluriclosed in action among a large class of metrics. This may for instance be used to establish the evolution behavior of many geometric quantities along these pluriclosed flow solutions.
Acknowledgements. We thank Beatrice Brienza and Ramiro Lafuente for fruitful conversations and to Fabio Podestà for bringing the paper [AD] to our attention.
@@ 2. Hermitian geometry of Lie groups
whenever $\alpha+\beta+\gamma=0$, where $y_{\alpha}=-\mathrm{i}\epsilon_{\alpha}x_{\alpha}$ and $[E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}$ (see §6), and
@@
=$=\left\{\begin{array}[]{l}0,\qquad\alpha+\beta\neq 0,\\ \omega_{\mathfrak{t}}(A,H_{\alpha}),\qquad\alpha+\beta=0.\end{array}\right.$
@@ 3. Pluriclosed metrics
A Hermitian structure $(J,g)$ is called pluriclosed if $dd^{c}\omega=0$ (or equivalently, ¯$\partial\overline{\partial}\omega=0$). We start this section by computing the $4$-form $dd^{c}\omega$ in our setting.
@@
Note that the above computation is not necessary if we use that $dd^{c}\omega=-\partial\overline{\partial}\omega$ is a $(2,2)$-form. On the other hand, for $\alpha,\beta\in\Delta^{+}$, $\gamma,\delta\in\Delta^{-}$, we have that
@@
$=$ $-2g_{\mathfrak{t}}(H_{\alpha},H_{\beta})-2N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})-2N_{\alpha,-\beta}^{2}\epsilon_{\alpha-\beta}(x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),$
@@
[…]
@@
[…]
@@
[…]
@@
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and for each $1\leq i\leq s$, there exists $z_{i}>0$ such that $g_{\mathfrak{t}}|_{\mathfrak{t}_{i}}=z_{i}g_{\operatorname{B}}|_{\mathfrak{t}_{i}}$ and $g_{\mathfrak{q}}|_{\mathfrak{q}_{i}}=(z_{i}x_{\alpha})_{\alpha\in\Delta^{+}_{i}}$, where the $x_{\alpha}$ are defined as follows: if $\Pi_{i}=\{\alpha_{1},\dots,\alpha_{n_{i}}\}$ are the simple roots of $\mathfrak{g}_{i}$ and we set $x_{j}:=x_{\alpha_{j}}$ for $j=1,\dots,n_{i}$, then
    $x_{\alpha}:=\sum_{j=1}^{n_{i}}k^{\alpha}_{j}x_{j}-\left(\sum_{j=1}^{n_{i}}k^{\alpha}_{j}-1\right),\qquad\forall\alpha=\sum_{j=1}^{n_{i}}k^{\alpha}_{j}\alpha_{j}\in\Delta^{+}_{i},\quad k^{\alpha}_{j}\in{\mathbb{Z}}_{\geq 0}.$
@@
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and on each simple factor, up to scaling, $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$ and $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ for all $\alpha,\beta\in\Delta^{+}$, where $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$.
@@
Given any of these metrics $g$, the space of all left-invariant complex structures which are compatible with $g$ depends on $d(d-1)$ parameters, where $2d=\operatorname{rk}(G)$. On the other hand, it follows from the theorem that for any fixed irreducible $J$, the space of all pluriclosed metrics compatible with $J$ depends on $12d+$ 1 parameters and contains a unique bi-invariant metric up to scaling, given by $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+\dots+z_{s}(-\operatorname{B}_{\mathfrak{g}_{s}})$. Note that $g_{\mathfrak{t}}=g_{b}|_{\mathfrak{t}}$ for any metric $g$ as in the theorem. In general, the number of parameters is $r(2d+1)$ and $J$ admits an $r$-parametric space of Hermitian bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold.
@@
We first prove that (i) implies (ii). Note that in the proof of Proposition 3.6, we also obtain assuming that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$ the following: $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ in cases (a) and (b), and $x_{2\alpha+\beta}=2x_{\alpha}+x_{\beta}-2$ in case (b). In case (c), condition (5) gives the following equations when applied to the pairs $(\alpha,\beta)$, $(\alpha+\beta,\alpha)$, $(2\alpha+\beta,\alpha)$ and $(3\alpha+\beta,\beta)$, respectively:
$-1=$ $x_{\alpha+\beta}-x_{\alpha}-x_{\beta},$
$-1=$ $4x_{2\alpha+\beta}-7x_{\alpha+\beta}+3x_{\beta}-x_{\alpha},$
$1=$ $5x_{3\alpha+\beta}-11x_{2\alpha+\beta}+6x_{\alpha+\beta}+x_{\alpha},$
$-1=$ $x_{3\alpha+2\beta}-x_{3\alpha+\beta}-x_{\beta}.$
This implies that $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$, $x_{2\alpha+\beta}=2x_{\alpha}+x_{\beta}-2$, $x_{3\alpha+\beta}=3x_{\alpha}+x_{\beta}-3$ and $x_{3\alpha+2\beta}=3x_{\alpha}+2x_{\beta}-4$. Summarizing, the formula in part (ii) holds for any two positive roots in $\operatorname{span}\{\alpha,\beta\}$, where $\alpha,\beta\in\Pi$.
We now proceed by induction in the height of the roots. A straightforward inspection of the nine types of Dynkin diagrams gives that for any $\gamma=\sum\limits_{i=1}^{2d}k_{i}\alpha_{i}\in\Delta^{+}$ such that at least three different integers $k_{i}$’s are $>0$, there exists $\alpha,\beta\in\Delta^{+}$ such that $\gamma=\alpha+\beta$ and $\alpha-\beta\notin\Delta$. Thus $N_{\alpha,\beta}^{2}=-\langle\alpha,\beta\rangle$ and so by (5), $x_{\gamma}=x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$. If $\alpha=\sum\limits_{i=1}^{2d}l_{i}\alpha_{i}$ and $\beta=\sum\limits_{i=1}^{2d}m_{i}\alpha_{i}$ (so $k_{i}=l_{i}+m_{i}$), then by induction hypothesis,
$x_{\gamma}=\sum_{i=1}^{2d}l_{i}x_{i}-(h(\alpha)-1)+\sum_{i=1}^{2d}m_{i}x_{i}-(h(\beta)-1)-1=\sum_{i=1}^{2d}k_{i}x_{i}-(h(\gamma)-1),$
as was to be shown. It is easy to check that part (iii) follows from part (ii).
Finally, if part (iii) holds, then we also have that $x_{\alpha-\beta}+1=\pm(x_{\alpha}-x_{\beta})$ if $\alpha-\beta\in\pm\Delta^{+}$ for all $\alpha,\beta\in\Delta^{+}$. Let us prove part (i) by using Corollary 3.4. We can assume that $G$ is simple and that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$, i.e., $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=\langle\alpha,\beta\rangle$ for all $\alpha,\beta\in\Delta^{+}$. By using that $N_{\alpha,-\beta}^{2}=N_{\alpha,\beta}^{2}+\langle\alpha,\beta\rangle$, we obtain that condition (5) for $\alpha,\beta\in\Delta^{+}$, $\alpha\neq\beta$ is equivalent to
@@
$=$ $N_{\alpha,\beta}^{2}(1+x_{\alpha-\beta}-(x_{\alpha-\beta}+1))+\langle\alpha,\beta\rangle(x_{\alpha-\beta}-(x_{\alpha-\beta}+1))=\langle\alpha,\beta\rangle,$
@@
If $\alpha+\gamma\in\Delta^{+}$, then by using that $\alpha=(\alpha+\gamma)+(-\gamma)$ and $-\delta=\beta+(-(\beta+\delta))$, we obtain
@@
Analogously, it is equal to $-2$ whenever $\alpha+\gamma\in\Delta^{-}$. On the other hand, if $\alpha+\delta\in\Delta^{+}$, then
@@
and it is equal to $-2$ if $\alpha+\delta\in\Delta^{-}$. This implies that (6) is equivalent to
@@ 5. Pluriclosed flow
where equality holds if and only if $\operatorname{grad}(F)_{X(t)}=0$ as $Q$ is positive definite, if and only if $X(t)\equiv X_{0}$. In other words, $F$ is a Lyapunov function for the dynamical system and it follows from standard methods that any solution $X(t)$ converges to $X_{0}$ as $t\to\infty$ (see e.g. [P, Section 2.9]).

### v3

@@ Abstract.
Given a compact semisimple Lie group $G$ and a maximal torus $T\subset G$, we give an explicit description of all left and $\operatorname{Ad}(T)$-invariant pluriclosed Hermitian structures on $G$ in terms of the corresponding root system. They depend on $2d+1$ parameters in the irreducible case, where $\dim{T}=2d$. As applications, we obtain that the only left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics which are also CYT are bi-invariant metrics (i.e., Bismut flat) and study the pluriclosed flow as a neat ODE system.
@@ 1. Introduction
Explicit examples of pluriclosed left-invariant metrics (non bi-invariant) were also exhibited in [Ph] for $G=\mathrm{G}_{2}$.
Remark 1.1.
It is worth recalling that given a pluriclosed metric $\omega$ on a compact manifold, $\omega+\overline{\partial}\psi+\partial\overline{\psi}$ is also a pluriclosed metric for any sufficiently small $(1,0)$-form $\psi$. Moreover, using that the $(1,1)$-Aeppli cohomology is one-dimensional (see [B, Theorem 3.1]), one obtains that this construction gives all left-invariant pluriclosed metrics on a compact simple Lie group.
@@
If $\operatorname{rk}(G)=2d$, then each of these metrics is compatible with a d(d1)d(d-1)-parametric space of left-invariant complex structures on $G$. According to the theorem, for a fixed $J$, if the space of all left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics compatible with $J$ is non-empty, then it depends on $r(2d+1)$ parameters and contains an $r$-parametric space of bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold. In particular, any irreducible $J$ admits at most one compatible bi-invariant metric up to scaling. Note that $J$ can be irreducible on a non-simple $G$ (see Example 2.1).
@@
and $F$ is the functional defined in (1). Thus $Y(t):=Q^{-\tfrac{1}{2}}X(t)$ is the negative gradient flow of the real analytic function $H(Y):=F(Q^{\tfrac{1}{2}}Y)$, which implies that any solution $X(t)$ converges to $(1,\dots,1)$ as $t\to\infty$. Any left and $\operatorname{Ad}(T)$-invariant pluriclosed solution $g(t)$ therefore converges, as $t\to\infty$, to a bi-invariant metric (see Theorem 5.1 for a more precise statement). We note that the ODE systems (2) and (3) clearly describe the pluriclosed flow evolution among a large class of metrics. This may for instance be used to establish the evolution behavior of many geometric quantities along these pluriclosed flow solutions.
Acknowledgements. We thank Beatrice Brienza and Ramiro Lafuente for fruitful conversations and to Fabio Podestà for bringing the paper [AD] to our attention. We are specially grateful to Joseph Kwong for pointing out a mistake in the proof of Theorem 3.7 in a previous version of this paper, and for helping us to fix it. We also thank the anonymous referee for Remark 1.1.
@@ 2. Hermitian geometry of Lie groups
whenever $\alpha+\beta+\gamma=0$, where $y_{\alpha}=-\mathrm{i}\epsilon_{\alpha}x_{\alpha}$ and $E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}$ (see §6), and
@@
$=\left\{\begin{array}[]{l}0,\qquad\alpha+\beta\neq 0,\\ \omega_{\mathfrak{t}}(A,H_{\alpha}),\qquad\alpha+\beta=0.\end{array}\right$.
@@ 3. Pluriclosed metrics
A Hermitian structure $(J,g)$ is called pluriclosed if $dd^{c}\omega=0$ (or equivalently, $\partial\overline{\partial}\omega=0$). We start this section by computing the $4$-form $dd^{c}\omega$ in our setting.
@@
Note that the above computation can be skipped by using that $dd^{c}\omega=-\partial\overline{\partial}\omega$ is a $(2,2)$-form. On the other hand, for $\alpha,\beta\in\Delta^{+}$, $\gamma,\delta\in\Delta^{-}$, we have that
@@
$=$ $-2g_{\mathfrak{t}}(H_{\alpha},H_{\beta})-2N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})-2N_{\alpha,-\beta}^{2}\epsilon_{\alpha-\beta}(\epsilon_{\alpha-\beta}x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),$
@@
    $\alpha$$\beta$ : $\langle\alpha,\alpha\rangle=\langle\beta,\beta\rangle=-2\langle\alpha,\beta\rangle$, $N_{\alpha,\beta}^{2}=\tfrac{1}{2}\langle\alpha,\alpha\rangle=N_{\alpha+\beta,\alpha}^{2}=N_{\alpha+\beta,\beta}^{2}$. It follows from (5) that
@@
    $\alpha$$\beta$ : $\langle\alpha,\alpha\rangle=\tfrac{1}{2}\langle\beta,\beta\rangle=-\langle\alpha,\beta\rangle$, $N_{\alpha,\beta}^{2}=\langle\alpha,\alpha\rangle=N_{\alpha+\beta,\alpha}^{2}=N_{2\alpha+\beta,\alpha+\beta}^{2}$. By (5) we have that $g_{\mathfrak{t}}(2\alpha+\beta,\beta)=0$, so $g_{\mathfrak{t}}(\alpha,\beta)=-\tfrac{1}{2}g_{\mathfrak{t}}(\beta,\beta)$, and in addition,
@@
    $\beta$$\alpha$ : $\langle\alpha,\alpha\rangle=\tfrac{1}{3}\langle\beta,\beta\rangle=-\tfrac{2}{3}\langle\alpha,\beta\rangle$, $N_{\alpha,\beta}^{2}=\tfrac{3}{2}\langle\alpha,\alpha\rangle=N_{2\alpha+\beta,\alpha}^{2}=N_{3\alpha+\beta,\beta}^{2}$, $N_{\alpha+\beta,\alpha}^{2}=2\langle\alpha,\alpha\rangle$. Here condition (5) gives that $g_{\mathfrak{t}}(2\alpha+\beta,\beta)=0$ and $g_{\mathfrak{t}}(3\alpha+\beta,\alpha+\beta)=0$, from which clearly follows that $g_{\mathfrak{t}}(\alpha,\beta)=\langle\alpha,\beta\rangle$ and $g_{\mathfrak{t}}(\beta,\beta)=\langle\beta,\beta\rangle$.
@@
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and on each simple factor, up to scaling, $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$ and $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$, where $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$.
@@
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and for each $1\leq i\leq s$, there exists $z_{i}>0$ such that $g_{\mathfrak{t}}|_{\mathfrak{t}_{i}}=z_{i}g_{\operatorname{B}}|_{\mathfrak{t}_{i}}$ and $g_{\mathfrak{q}}|_{\mathfrak{q}_{i}}=(z_{i}x_{\alpha})_{\alpha\in\Delta^{+}_{i}}$, where the $x_{\alpha}$ are defined as follows: if $\Pi_{i}=\{\alpha_{1},\dots,\alpha_{n_{i}}\}$ are the simple roots of $\mathfrak{g}_{i}$ and we set $x_{j}:=x_{\alpha_{j}}$ for $j=1,\dots,n_{i}$, then x_α:=_j=1^n_i k^α_jx_j -(_j=1^n_i k^α_j -1),   α=_j=1^n_i k^α_jα_jΔ^+_i,  k^α_jZ_0.
@@
Given any of these metrics $g$, the space of all left-invariant complex structures which are compatible with $g$ depends on d(d1)d(d-1) parameters, where $2d=\operatorname{rk}(G)$. On the other hand, it follows from the theorem that for any fixed irreducible $J$, the space of all pluriclosed metrics compatible with $J$, if non-empty, depends on $2d+1$ parameters and contains a unique bi-invariant metric up to scaling, given by $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+\dots+z_{s}(-\operatorname{B}_{\mathfrak{g}_{s}})$. Note that $g_{\mathfrak{t}}=g_{b}|_{\mathfrak{t}}$ for any metric $g$ as in the theorem. In general, the number of parameters is $r(2d+1)$ and $J$ admits an $r$-parametric space of Hermitian bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold.
@@
We first prove that (i) implies (ii) by induction in the height of the roots. On any of the simple factors of $G$, we can assume by Proposition 3.6 that $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=\langle\alpha,\beta\rangle$ for all $\alpha,\beta\in\Delta$, so it follows from (5) that
$\langle\alpha,\beta\rangle=N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})+\epsilon_{\alpha-\beta}N_{\alpha,-\beta}^{2}(\epsilon_{\alpha-\beta}x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),$
for all $\alpha,\beta\in\Delta^{+}$, $\alpha\neq\beta$. If $\epsilon_{\alpha-\beta}=0$, then $N_{\alpha,\beta}^{2}=-\langle\alpha,\beta\rangle$ and we are done. Otherwise, we can assume that $\epsilon_{\alpha-\beta}=1$, i.e., $\alpha-\beta\in\Delta^{+}$, and since $x_{\alpha}=x_{\alpha-\beta}+x_{\beta}-1$ by induction hypothesis, we obtain that (see §6)
$\langle\alpha,\beta\rangle=N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})+N_{\alpha,\beta}^{2}+\langle\alpha,\beta\rangle,$
so $x_{\alpha+\beta}-x_{\alpha}-x_{\beta}=-1$.
On the other hand, it is easy to check that part (iii) is equivalent to part (ii).
Finally, if part (ii) holds, then we also have that $x_{\alpha-\beta}-1=\pm(x_{\alpha}-x_{\beta})$ if $\alpha-\beta\in\pm\Delta^{+}$ for all $\alpha,\beta\in\Delta^{+}$. Let us prove part (i) by using Corollary 3.4. We can assume that $G$ is simple and that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$, i.e., $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=\langle\alpha,\beta\rangle$ for all $\alpha,\beta\in\Delta^{+}$. By using that $N_{\alpha,-\beta}^{2}=N_{\alpha,\beta}^{2}+\langle\alpha,\beta\rangle$, we obtain that condition (5) for $\alpha,\beta\in\Delta^{+}$, $\alpha\neq\beta$ is equivalent to
@@
$=$ $N_{\alpha,\beta}^{2}(-1+x_{\alpha-\beta}-(x_{\alpha-\beta}-1))+\langle\alpha,\beta\rangle(x_{\alpha-\beta}-(x_{\alpha-\beta}-1))=\langle\alpha,\beta\rangle,$
@@
If $\alpha+\gamma\in\Delta^{+}$, then by using that $\alpha=(\alpha+\gamma)+(-\gamma)$ and $\delta=\beta+(-(\beta+\delta))$, we obtain
@@
Analogously, it is equal to 2-2 whenever $\alpha+\gamma\in\Delta^{-}$. On the other hand, if $\alpha+\delta\in\Delta^{+}$, then
@@
and it is equal to 2-2 if $\alpha+\delta\in\Delta^{-}$. This implies that (6) is equivalent to
@@
Remark 3.10.
The above proof shows that condition (6) actually follows from (5) in the characterization of pluriclosed given in Corollary 3.4.
@@ 5. Pluriclosed flow
where equality holds if and only if $\operatorname{grad}(F)_{X(t)}=0$ as $Q$ is positive definite, if and only if $X(t)\equiv X_{0}$. In other words, $F$ is a Liapunov function for the dynamical system and it follows from standard methods that $X_{0}$ is asymptotically stable, in the sense that any solution $X(t)$ starting near $X_{0}$ converges to $X_{0}$ as $t\to\infty$ (see e.g. [P, Section 2.9]). On the other hand, if we change to the variable $Y:=Q^{-\tfrac{1}{2}}X$ and consider the new functional $H(Y):=F(Q^{\tfrac{1}{2}}Y)$, then it is easy to check that $X(t)$ is a solution to (14) if and only if $Y(t)=Q^{-\tfrac{1}{2}}X(t)$ is the negative gradient flow of $H$:
$Y^{\prime}(t)=-\operatorname{grad}(H)_{Y(t)}.$
Since $Y_{0}:=Q^{-\tfrac{1}{2}}X_{0}$ is a global minimum and only critical point of $H$ and the function $H$ is real analytic, it is well known that indeed any solution $Y(t)$ converges to $Y_{0}$ as $t\to\infty$.